OpenStudy (momwasright):

Given the functions f(n) = 25 and g(n) = 3(n – 1), combine them to create an arithmetic sequence, an, and solve for the 12th term

OpenStudy (momwasright):

an = 25 – 3(n – 1); a_12 = –11 an = 25 – 3(n – 1); a_12 = –8 an = 25 + 3(n – 1); a_12 = 58 an = 25 + 3(n – 1); a_12 = 61

OpenStudy (momwasright):

nvm i got it right thanks :)